OpenStudy (anonymous):

Find dy/dx if y=1+sinx/1-sinx

OpenStudy (anonymous):

\[\frac{ (cosx)(1-sinx)-(1+sinx)(-cosx) }{ (1-sinx)^{2} }\]by the quotient rule

OpenStudy (anonymous):

which simplifies to \[\frac{ 2cosx }{ (1-sinx)^{2} }\]

OpenStudy (arnavguddu):

i guess it will be (1+sinx)/(1-sinx) then derv = ((cos x)(1-sinx) - (-cos x)(1+sin x))/(1-sin x)^2 ...